Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-02-10
ISS On-Orbit Status 02/10/08

Sunday --- Mission 1E Flight Day 4 (FD4).

Ahead: Week 17 of Increment 16.

Mission 1E replanning by MCC-Houston, driven by the one-day delay of the first spacewalk, was completed last night, as follows:

Approved mission extension by one day (i.e., 12+0+2 instead of 11+0+2), resulting in 2/19 (Tuesday) as return date for Atlantis;
Limiting resource being oxygen (O2), under further discussion is the option to either extend by one more docked day or transfer the unexpended O2 to the ISS (current O2 margins are estimated to be 11+1+2 plus an additional 20-25 hours. These are still being refined given the insertion of the new FD4);
All FD 4 activities moved to FD 5. EVA-1 will be conducted by Stanley Love and Rex Walheim tomorrow, with Campout of the two tonight;
Today (FD 4) was replanned and includes 1.5 hrs of focused inspection (FI) of the starboard OMS (Orbital Maneuvering System) pod blanket that has a corner slightly peeled back (the FI, starting at ~ 2:15 pm EST, is planned for 90 min, with additional FI time available on FD 06 if needed);
Transfer status: 22% of transfer complete, 6 hours of transfer scheduled for FD 4, expect 35/40 lbs N2 transfer today, 3 CWCs filled.
Wakeup time for the ISS crew remains at 4:45am EST, with sleep time tonight 8:45pm. Same times for the Shuttle crew.

After wakeup and before breakfast, FE-2 Dan Tani completed his daily access of the SLEEP experiment (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) software for data logging and filling in questionnaire entries in the experiment's laptop session file on the HRF-1 laptop for downlink. (To monitor the crewmember's sleep/wake patterns and light exposure, Dan wears a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

In Node-2, CDR Whitson performed a leak check on the O2 supply line set up yesterday to allow flow of Shuttle O2 to the ISS PBAs (Portable Breathing Apparatus) in support of pre-EVA mask prebreathe for denitrogenation.

The crew, particularly CDR Frick, PLT Poindexter, MS2 Walheim and MS4 Love, conducted an in-depth review of procedures for the EVA-1 spacewalk, with egress scheduled to start tomorrow morning at ~9:40am EST.

CDR Whitson, with Frick and Poindexter, initiated N2 (nitrogen) transfer from the Orbiter to the ISS HPGTs (High-Pressure Gas Tanks) on the exterior of the U.S. Airlock (A/L). (A total of ~40 lbs of N2 was to be transferred to the ISS (until stopped by the HPGT counter pressure).)

FE-1 Malenchenko transferred and replaced new Russian ODF (Operations Data Files) documents and updates from Progress 28P. (The new material was inserted in five Procedures Books (SOZh Life Support System, SOGS Atmosphere Revitalization System, VKS Auxiliary Computer System, PTO VnuKD FGB IVA IFM (In-flight Maintenance) Part 1, and RPR TKG #363 Progress M-63 Cargo Transfer Ops.)

Similarly, CDR Peggy Whitson transferred a CD (Compact Disk) storage box with new data files for the 1E stage from Atlantis to the ISS.

Malenchenko completed the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the Service Module (SM). (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.)

Yuri also handled the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

More CWC (Contingency Water Container) bags were filled in the Atlantis with water from the Shuttle's fuel cells by Frick & Poindexter for transfer to ISS.

Yuri Malenchenko performed a successful leak check on his Russian Sokol spacesuit.

More 1E cargo transfers from the Shuttle to ISS were handled by Poindexter, Frick and Whitson.

FE-2 Dan Tani, slated for return on the Atlantis on 2/19, and MS5 LĂ(c)opold (Leo) Eyharts, who is to replace him as ISS FE-2, had 3h 15m scheduled for standard joint handover activities, to be continued through the docked period ahead.

Tani downlinked the data from his two CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS) Actiwatches and body-worn Holter recorder.

EV1 Walheim & EV2 Love worked on their EMUs (Extravehicular Mobility Units), adjusting sizes and swapping components as required (MS3 Hans Schlegel being scheduled for spacewalk next week). Preparations for tomorrow's spacewalk included refilling the in-EMU DIDBs (Disposable In-suit Drink Bags). (The DIDBs were filled yesterday with non-iodinated drinking water, good for 24 hr & intended for today. The postponement of EVA-1 to FD5 required a refill with fresh non-iodinated potable water.)

Schlegel meanwhile checked out the REBA (Rechargeable EVA Battery Assembly)-powered EVA hardware in the Airlock (A/L), as Rex and Stan prepared the A/L Equipment Lock for their campout. (Reminder: The U.S. A/L has two hatch-separated compartments: the Crew Lock (CL) and the Equipment Lock (EL).)

In Node-2, Peggy prepared PCS (Portable Computer System) 120 Vdc UOP (Utility Outlet Panel) power for the spacewalk and verified correct A31p laptop function.

Whitson also started the outfitting of the Node-2 port vestibule for tomorrow's berthing of Columbus (COL). The vestibule outfitting will be continued over the next several days after the COL berthing.

Schlegel, Love & Tani each have a PFC (Private Family Conferences) on their schedule, via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on the SSC-9 laptop), Stan at ~1:50pm, Hans at ~5:30pm, Dan at ~6:15pm.

At ~2:15pm EST, Shuttle crewmembers Leland Melvin (MS1), Frick, Poindexter and Love began the 2.5-hr. focused inspection of the starboard OMS pod with the detached corner of one of the TPS (Thermal Protection System) blankets. The resulting imagery will be downlinked later tonight for inspection by ground specialists. (The forward corner of the blanket (~6 in x ~6 in) is peeled back and sticking up approximately 2 inches. There is no RTV visible and there is no structural rib underneath. Additionally, there is a loose portion of the blanket in the middle raised to an unknown height.)

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR), TVIS treadmill (FE-1, FE-2), RED resistive exercise device (CDR, FE-2), and VELO bike with bungee cord load trainer (FE-1).

Yuri transferred the crew's exercise data file to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Working off his discretionary 'time permitting' task list, Malenchenko conducted the daily monitoring, picture-taking and downloading on the newly set up BIO-5 Rasteniya-2 ("Plants-2") experiment. (Rasteniya-2, researches growth and development of plants (peas) under spaceflight conditions in the Lada-12 greenhouse from IBMP (Institute of Bio-Medical Problems, Russian: IMBP). During its operation, the experiment requires regular daily maintenance of the experiment involving monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, and photo/video recording. During the duration of the BIO-5 experiment, students of the Moscow City Palace for Youth Creativity of the Meshchansky inter-regional center #15 in Moscow) and the Prince of Oldenburg Lyceum in St. Petersburg will be cultivating plants in parallel on the ground and conducting comparative observation of plant growth and development under gravity and zero-gravity conditions. They are receiving the photo images taken by Yuri.)

Rex (EV1) and Stan (EV2) will begin their 'campout' (nachalo desaturatsiy = desaturation start) in the A/L with hatch closure and depressurization of the CL from 14.7 to 10.2 psi at ~7:00pm, followed by mask prebreathe at ~7:00-8:15pm. Sleep time for the ISS crew begins at 8:45pm. (For the Campout, METOX (Metal Oxide) canisters #0017 & #0019) have been installed in the A/L for CO2 control.)

Following the usual hygiene break/with mask prebreathe for Walheim & Love at ~5:20-6:30am tomorrow morning after spending the night on 10.2 psi, the A/L hatch will be closed again for EVA preps in 10.2 psi, followed by EMU purge and prebreathe in the EMUs. Afterwards, with CL depressurization and EV1/EV2 egress, EVA-1 nominally begins at ~9:40am EST. Frick & Poindexter will support the spacewalk as IV (Intravehicular) crewmembers, keeping tabs with the detailed activity steps and crib sheet.

Unberthing and transfer of the COL Module to its Node-2 CBM (Common Berthing Mechanism) by Dan Tani, Leo Eyharts and Leland Melvin will begin tomorrow at ~2:10pm.

No CEO photo targets uplinked for today.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use